Find g(x),the inverse function of f(x)=3x/x+1.
The correct answer is one the following. Which one is correct?
A) g(x)=x/3+x
B) g(x)=x/3-x
C) g(x)=3-x/x
D) g(x)=x-3/x

Answer by Edwin McCravy(20081)

When you use a slash " / " for a fraction bar, you must use parentheses 
before and after it to show what is the numerator and what is the denominator.

Did you mean this:

Find g(x),the inverse function of f(x)=(3x)/(x+1).
A) g(x)=(x)/(3+x)
B) g(x)=(x)/(3-x)
C) g(x)=(3-x)/(x)
D) g(x)=(x-3)/(x)

Or did you mean this:

Find g(x),the inverse function of f(x)=(3x)/(x)+1.
A) g(x)=(x)/(3)+x
B) g(x)=(x)/(3)-x
C) g(x)=3-(x)/(x)
D) g(x)=x-(3)/(x)

Please post again using parentheses to show what goes in the 
numerator and denominator and what doesn't.

Edwin